OpenStudy (anonymous):

The eight grade class received surveys regarding Internet access in their homes. Three-fourths of the eight grade class has access to the Internet in their homes. If 120 students have access to the Internet, how many students are in the eighth grade class? Explain.

OpenStudy (anonymous):

160

OpenStudy (anonymous):

3/4 = 120 1/4 = 40 ( 120 / 3) 4/4 = 160